3. The maximum wavelength for photoelectric emission in tungsten is 230 nm. What wavelength of light must be used inorder for electrons with a maximum energy of 1.5 eV to be ejected ?

Please solve in detailed steps, thank you
Answer #1

Here ,

let the wavelength of the light to be used is L

Now, as the cut off wavelength , L0 = 230 nm

total energy of the photon = cutoff energy + kinetic energy

6.626 *10^-34 * 3 *10^8/(L * 1.602 *10^-19) = 6.626 *10^-34 * 3 *10^8/(230 *10^-9 * 1.602 *10^-19) + 1.5

NOw , solving for L

L = 1.8 *10^-7 m = 180 nm

the wavelength needed is 180 nm